A vessel has 4 litres and 500 ml of curd. In how many glasses, each of 25 ml capacity, can it be filled?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Quantity of curd in the vessel = 4 l 500 ml = 4×1000 + 500 = 4500 ml
Capacity of 1 glass = 25 ml
∴Number of glasses that can be filled with curd = 4500 / 25 = 180 glasses Hence, 180 glasses can be filled with curd

Thanks